OpenStudy (anonymous):

The y-intercept for 3x + 2y = 12 is (0,___). A. 3 B. 2 C. 6 D. 4

OpenStudy (mysesshou):

Agree! 2y = -3x + 12 y = -3/2x + 6 b=6 intercept = 6